by TEI TEI

03/12/08 – TTC modified the final punchlist again and sent it out to JCI and Triton along with the final as-built drawings TEI had on the project. TEI email Banta to see about the final as-built drawings that were received on this project from Banta on this job. Punchlist 3 & 4 uploaded to the web.

12/19/07 – TEI called for an owner witnessed fire alarm test for MJ. Banta said they could provide this. JCI won’t pick up the fire alarm but Facilities can see how the unit will go into alarm when the smoke pulls out. John Sanders he will work on finishing all the work on this project and stop by today. Boiler 2 had a loose ignition wire and Ken Morris fixed that last Friday but will send the service ticket to the County.

12/17/07 – Triton and JCI were on site last week trying to closeout the project. Today MJ reported that smoke detectors appeared jumpered out on the system and one was missing. Banta said they know this not to be the case when they tested them and will send Ed Ginter over to verify operation of the smoke detectors. TTC forwarded the Dowtherm white paper to MJ & KM showing the 30% provides burst protection down to -31degreeF so 30% ethylene glycol should be okay.

12/12/07 – JCI was onsite today to provide final training on the DDC systems on the project. TTE showed John Sanders some items that were not complete on the project. The freezestats were un jumpered but now the units are going to have trouble running through the cold weather because of the freezing cold air coming into the unit in the economizer mode. JCI reported that the freezestats cannot be set below the 35F factory setting but TTC will check to see about another setting or different equipment. The glycol mix did not prove to be protected down to -20F nor does the concentration seem to be at 30%. The Chardon lab chart says 30% will get 5F for this system. TTE to reconcile.

12/04/07 – TTC has been working with Triton & JCI to get the project completed and commissioned. TTC told KM today that he believes all the contract work is complete for all contractors and that we can move to just warranty issues. TTC asked JCI to schedule final training on DDC system for this Friday but this date is not confirmed. TTC has scheduled the final back punchlist visit for this Thursday. TTC contacted Banta about missing information in the O&M manuals such as warranty letter, contact info and all closeout documents. KM asked for final Banta payment today and TTC called Banta and told them to send both final payapps to TEI today for approval. TTC subsequently spoke to Banta about final change order cost (3 changes listed below) and update the website. Banta appears ready to send over the payapps for this job. TTC reported all of this to KM. MJ has asked several times about the glycol % concentration in the loop and TTC has sent he spec required 30% back to KM each time. This is what TEI told Triton to buy and they provided premixed glycol and put it into the system. Banta sent the payapps over and TTC signed them and sent them via City Dash down to KM at the CH.

11/16/07 – TTC issued the second mechanical and control punchlist today to Grady and Brett for immediate completion. PDF

11/13/07 – Tim Centers called today to report that Habegger will be on site tomorrow to check the toggle position of the incoming voltage. If we set it to 208V he believes that a lot of our over starts will go away and our long wait time in the summer for chiller restarts on power loss will also be very diminished. He was still willing to meet Facilities at the site to discuss the chiller programming next week. KM sent in many questions about the chiller plant and programming and TTC forwarded these to JCI for answers. Most of the final details lie in the alarm programs which JCI is willing to extend to 5 minutes to reduce alarm reports on the pumps, chiller plant and boiler plant. TEI has asked that low level alarms be removed from AHU-3 because it is scheduled off almost all the time.

11/02/07 – TEI sent the Engr Change Order to Kevin for finalization. MJ emailed today to report the compressor are banging on/off way too much. TTC has not witnessed this under the latest sequence and the timing seems to be about once every 15 minutes in the field. We will however bring Carrier over in 2 weeks to review with owner.

11/01//07 – Triton finished up all the paperwork and delivered it to TTC at TEI. TEI sent to Kevin via City Dash. The O&M’s were delivered but Grady and TTC will send to County for testing. All final inspections have been passed for both Building, HVAC and Seismic.

10/17/017 – TTC sent out a large email outlining the items left at the CC. Grady reported that final balance reports will be here by Friday. Some mechanical work remains like OA boots, insulation and inspections but the project is almost complete. It is just programming that needs to be ironed out at this point.

10/01/07 – TTC complete 2.5 days of commissioning at the Comm Center. JCI was not ready for CX and TTC could not back check the reports they filled out. We worked out many issues with the boilers and air handlers but still have more issues to resolve. TTC asked CC about leaving the ladder and door arrangement as it is right now after the ladder was moved. TTC contacted Triton and JCI about finishing the project ASAP. TTC requested trend reports for verification of working systems.

09/24/07 – Kevin reported the room was overheating much of Friday and JCI was on site today attempting to speed up the sequences. Kevin said it was 73 F which I don’t consider overheated but it will shake out in the commissioning reports. TEI is scheduled to back commission the project starting Tuesday afternoon. TTC approved and sent out Triton payapp 5. Cousins inquired about their payapp 3 which TEI said they sent to PW in August. Banta has never billed us. TTC asked for the new receptacle for the water softener again.

09/19/07 – TTC sketched up the new control scheme for Return Fan and Relief Fan operation for AHU-2. TEI sent the sketches to JCI, Triton and the County for implementation into the project.

09/12/07 – TTC performed the mechanical punchlist yesterday. Most equipment looks as though it is in very good shape. The insulation on AHU-2 looks fantastic. TTC wants to submit the punchlist to Triton today.

08/28/07 – Training is schedule for tomorrow. The boiler boards did not arrive yet. Denny said maybe tomorrow. Airflow stations are not in yet but they are being ordered. Triton updated their work to do list and TTC sent it to KM and TM. Triton asked for written confirmation on moving the ladder in the storage room.

08/23/07 – KM stopped by this morning and reported that space temps were 72-75 but humidity was up in both areas. Trend data shows humidity rises during shift changes. JCI said they are providing a laptop for interface into the JCI system on site and we discussed agenda items for training. A new boiler control panel board must be ordered to give the boilers a firing rate contrary to what Denny Lammers told me on the phone yesterday. The new boards will be here on Monday and installed immediately by Scott Black. Denny Lammers called to say that the comm ports on the boards in the boilers at the Comm Center have been shorted out and new ones will have to be ordered. The should be here today or tomorrow.

08/22/07 – KM stopped by this morning and humidity was up at 9am but came right back down after the doors were closed and the shift change was finished. KM asked for a trend point to track the humidity to see when it is rising up so high (66% in Admin area). The 911 center was fine but Greg said the temp was 72 even though the stat was set at 68. These are averaging stats and this makes sense. Each stat can be overridden with its control knob. Bill thought his office was a little stuffy but his temp was 72 while the other office was 68 in AHU-1 zone. We could change the setpoint of this space down to 70. TEI asked for a copy of current sequences to be given to the owner. Training has been scheduled for next Tuesday at 8am. Denny Lammers confirmed that the boiler firing rate can be fed directly to the boiler with a 0-10V signal from the JCI system. The DDC system will also select the lead boiler and let the County manually rotate the boilers. JCI said that graphics might be installed by today on the Comm Center computer.

08/21/07 – KM called after he stopped out on site. Temps are good. JCI has everything programmed in automatic at this point except the boiler control which is being finalized and worked on with Denny Lammers. Greg said everything was delightful and Triton was working to finish everything up.

08/20/07 – TTC stopped at the site over the weekend to adjust temperature controls because the space was too hot and too cold and humidity was up. TTC sent an email last Sunday outlining all that was done over the weekend to make the systems work in a more automatic control.

I am writing this to outline what all happened at the Comm Center over the weekend and since last Thursday when JCI uploaded the new programs. You may need some of this information to release overrides when you program the system in its totality on Monday (tomorrow).

We have had a lot of overcooling since the new programs were put into place last Thursday and the only real control that I know we have is to manually adjust the chilled water valve to compensate for space temperature. Obviously this is not a good solution because we either overcool overnight or undercool in the morning. There is no one setting to keep the building temperate all the time.

On Friday, Kevin Maue went to the site and made some changes because the main area (AHU-2) was overcooling. With a little investigation, I believe we found that the supply air reset temperature was again programmed backwards meaning that as the space got cooler the discharge air temperature got lower. Mike Jackson noted this on the original JCI control submittals and it was caught again by John at JCI who went into the program and changed the supply air reset schedule when he was there one day according to a voicemail from Tim Myers to me. The headings on the JCI sequence of operation are backwards and the space temperature numbers are labeled as discharge air temperatures of the unit and vice versa. I believe when the new program was uploaded we went back to the old wrong way because the space was 64 degrees on Friday and the cooling valve was open 100% the humidity was below 55% so I don’t believe the unit thought it was in the dehumidification cycle. I had Kevin back the cooling valve down to 25% for the day. In the short term, this works but not overnight when it cools down.

On Saturday when Tom Rice from Triton was on site the space temperature rose to 76 degrees and the humidity was high. I stopped in about 1:30 p.m. and opened the cooling valve to 50% and drove the space temperature back down before I left. I also checked the supply air fan and return air fan speed which were okay but were manually set at 40 Hz and 30 Hz. I told and showed Tom how to adjust the cooling valve back to 25% before he left to prevent overnight overcooling (I believe he forgot to do this).

On Sunday, Mike Jackson called me at 9:30 a.m. to say it was 60 degrees in the space. I got the message at 4pm and drove to the Comm Center and arrived about 4:30 p.m. They reported that the coldest it got was 64 degrees and I found the chilled water valve still open at 50% which explains why it was so cold. I believe I was the only one to come out on Sunday to address this problem. Figuring this isn’t going to work overnight again, I decided to see if I could get the boiler system running on manual while I continue to wait for the final controls to be installed. On the boiler control panel it appeared to be locked out on outside air temperature. I got the user manual off the table, entered the passcode, browsed to the OAT disable point and raised it to 100 degrees F. I went into the AHU-1 room and put the hot water pump in hand. I check all the HW valves and they appeared to be open. I went to the front end and commanded the boiler on. It went into alarm. I walked back to the boiler and saw the LCD said LWCO lockout. I checked the new recirculation pump and it was off. I went back to front end and commanded the boiler pump # 1 on. I went back to the boiler and hit the reset switch on the boiler and it post purged, pre purged, ignited and started making hot water. I check the AHU-2 conditions and the space was still too cold and the humidity didn’t seem to be getting any lower either. I commanded the OA dampers closed and I finally commanded the cooling valve to 90%. After that I saw the heating valve start to crack open so I figured it was responding automatically to the space conditions. After about 45 minutes, the AHU-2 unit had 72 degree space conditions and 51% RH AND IT SEEMED TO BE OPERATING IN A SOMEWHAT AUTOMATIC SEQUENCE because the hot water valve kept stroking based on space conditions. I am hopeful that this will prevent overcooling/overheating until tomorrow when our contractors will be back to finish up.

I couldn’t get the hot water valve to open automatically on AHU-1 so I left the CHW valve at 20% chilled water position. The SF VFD is not reporting any speed and the VFD is actually in the hand position at 35%. The RELIEF fan is still OFF the way it should be. I closed the OA dampers on this unit too for now.

The AHU-3 unit cycles on about every 10 minutes. I don’t think that is necessary in the final program but we will see if it acts the same when we are finished programming it.

One thing is for sure, I am not comfortable, nor do I believe, that I should be the one making these changes in the field to this equipment. We, as a team, have to get this job completed immediately. What is it going to take to wrap up all the controls except for the air flow stations? Please do that immediately. If Denny Lammers has to come out to reprogram the boilers and coordinate with JCI, why hasn’t he been out there yet? We have given him work and spec’d his boilers on jobs in the past year. At the MRDD I accepted three of these boilers this summer. Please tell him that I expect him to be at the Comm Center on Monday with JCI getting the heating system running in automatic mode per our latest sequence. WE have to get this system running in automatic mode and eliminate these wide temperature swings. The 911 center is a 24/7 operation. Please keep coming out each day until it is finished.

Here are other comments that I have from the weekend:

Triton worked on Saturday and I saw Tom had the new water line routed from the basement to the upstairs bathroom for the humidifier. I don’t know if he got that finished as I left before he was done.
The water softener was not there. When is this coming in?
The insulation on AHU-2 looks worse then even. Now there is a big air pocket bubble where the ductwork isn’t sealed under the insulation and the air must be pushing the insulation out. This is near the top elbow going into the building. The insulation job on this unit is not acceptable. Rip it all off and replace it with the same type that I have the MRDD building. Please have them do a neat professional job when they replace this stuff so there will be no question about accepting it on the punchlist.
The supply air temperature on AHU-1 has a minimum setpoint of 55 degree F on one of the variable near the bottom of the JCI points list but the DAT was 53 degrees today. How can that be?
The AHU-1 VFD speed says 0% but the unit is running at 35 Hz in the Hand position. Is this a status point or a signal we are feeding the unit? If the unit is occupied as it said it was why is the SF speed at 0? I could not figure that out.
Why is the OA temp unreliable? Why is it reporting different temperatures on AHU-1 and AHU-2? One said 75 and one said 95. Don’t we have a global OAT sensor?
The hose bibbs are complete. One was rebuilt, the other is new.
How come we don’t have manual overrides at the JCI control valves for all coils? I have to check again but I am almost positive that GD always gives us this option in his specs. I could not find anywhere on these valves that I could do that.
When are we going to get the t-stats with overrides in the space? Maybe they are there – if so, just let me know. It would be a way for day shift/night shift to make small adjustments to setpoint.
When are the new diffusers going into the Radio Room area?
When are the going to insulate the remaining basement ductwork?
When is Tom going to cap off the abandoned ductwork? Reroute the existing?
Any word on why the chiller seems to take 15 minutes to restart on power failure? Tom said the other day he was just scrolling through the menus when it came back on full time. That is kind of what happened the other day too.

And finally, Tim I was glad to hear that the premix glycol is here. Please give me the change order cost to put the thermometers on each side of the coils like we discussed in my email. I want that work done ASAP as well and do it when the system is down for glycol. Troubleshooting will be a lot easier when those thermometers are installed. We don’t want to commission anything until the system has glycol in it because the system efficiency changes.

08/16/07 – Duke had another power outage at the Comm Center today. Triton & JCI were onsite. Tom Rice did not about the chiller not having to be locked out during generator operation any more. MJ contacted TTC and TTC emailed MJ and KM back about the situation explaining Tom just hasn’t read the meeting minutes from our control meeting.

08/13/07 – Today the spaces are comfortable. Triton came in over the weekend to work on the relief damper for the AHU-1 relief fan. They turned and left off the return fan to the unit. They could not confirm when JCI was going to finish their work on this job. TTC & MJ turned the fan back on and slowed both the supply and return back down to 35% and 25%. TTC issued meeting #12 today to Triton and the Owner. PDF

08/08/07 – MJ reported some cooling issues in the Admin wing. JCI and Triton check onsite for problems. TTC sent the following email to surmise the happenings:

I got a call yesterday that the Admin offices are warm at the Comm Center. Mike Jackson reported that the RF was shutdown, the cooling valve was only at 20% open and we are not sure how fast the SF is running. MJ stopped by for me and put the RF back into 50% speed, opened the CHW valve to about 50% and the space started to cool down. I called JCI and dispatched a service tech to the site this morning to check over the controls and they reported that the space and the trend data showed the space setpoint temp was set to 73 degree F and that the AHU was keeping that temperature okay. We had them reset the unit down to 70 degree F before they left. The space still seems warm today by local accounts. I called the comm center and talked to Lisa this morning who said the general feeling was that there was low air and it was stuff and warm. I called JCI and Triton and they agreed to send a JCI tech out to the building again to check the status of the unit and call me back. My goal is to increase the airflow and cool the space down a bit more. I also want to see what the actual space temperature is independent of the temperature sensors (MJ said he might be able to put a chart recorder in the space for me). I also asked JCI what the status of the existing controls are and although they could not say for sure 100% but they did say that the system is automatically responsive to the temperature sensors in the space. Based on the TEI sequence if the space is really overheating, both the VFD would normally be indexed to a higher speed and the SA discharge temp on the unit would be decreased. We do not have automatic control of the VFD yet (that happens next week) but we do have control over the CHW valve. The only reason it is not opening more is that the averaging t-stats in the admin wing are saying that they are meeting their setpoint. The chart recorder to tell us different or I may need to have JCI relocate one of the sensor to more representative are. We will know a little more in a bit and by the end of August we should be under 100% control with everything working automatically.
08/06/07 – Banta reported today that they have finished their punchlist including gate bonding. All documents are being sent to TEI for approval and closeout. Triton continues to work on the airflow monitoring stations for this project.

08/02/07 – Due to another power outage, JCI has to reset AHU-3 again at the Comm Center. They will be out there today to do this work. TTC & KM talked about completing this work ASAP. TTC sent out Electrical Punchlist today. PDF

07/19/07 – TTC called Cousins about Payapp #3 and #4.

0718/07 – AHU-3 had a power outage and would not stay running last week. Banta was able to provide a new conduit and feeder to the unit to fix the problem. Triton gave us a commitment of balancing to being on July 30 and TEI schedule a design meeting on July 25 for controls and owner training. Triton still owes TEI and Kevin Tom’s Work to Do List for Comm Center.

07/03/07 – TTC schedule the DDC meeting for July 12th for everyone to attend. TTC spoke to the water softener vendor about the specs on our new water softener in the change order for the project. Greg Wenz had some questions and comments about the system and power outages which are a problem now but will be taken care of upon completion of the DDC system. TTC reiterated to the County that JCI fully meets the spec and that the equipment is actually an upgrade to what Mike Jackson asked about being in the base bid. There are still some issue to resolve on this project. TTC owes Triton a sketch for the water softener installation and the new insulation scheme for the ductwork in the basement. We are not sure how far the balancers are don with their job so far and we need a decision on how we are replacing the diffusers in the ceiling of the 911 center.

06/26/07 – Tim is on vacation. We are trying to get the project wrapped up. Balancers will be on site tomorrow. Tom Rice is on site doing finishing work for the project. TTC contacted JCI to try to force them to get the project finished on the control side. TTC has a boiler RFI to answer for JCI at both buildings.

06/07/07 – Triton is scheduling the balancers for the main building. TTC is scheduling the punchlist for Friday at 2pm. TTC found payapp #2 from Triton and is going to hand deliver tomorrow morning.

06/04/07 – ttc approved triton’s payapp #4 for comm center. tom rice reported that chiller flow switch is now installed on chiller. fencing is installed but not the netting. gates are up around the equipment. ttc visited site on Saturday morning to check on progress from parking lot.

05/31/07 – All units are up and running. HHW piping is being completed this week. Duct cleaning is finished. Roof exhaust, relief and powered exhauster are installed but not finished. Insulation is not complete. Triton owes me a revised price on minimal ductwork insulation for old HV ductwork. We have not resolved the new diffuser install issue with round necks vs square necks but Tom and Tim are working on that. TEI plans on doing their initial punchlist early next week on site. The fencing is installed and the fabric is getting set to be installed. Space conditions seem to be quite well for the outdoor air temp. TEI issued meeting notes for Const Mtg #11. The next mtg is TBD. Triton submitted payapp #3 and ttc approved and sent to km.

05/17/07 – AHU-1 was shutdown and replaced and turned on today at 1:00 pm. Duct cleaning is schedule to be completed tonight at 10pm. TTC setup a Const Mtg next Wed at 2:30 p.m. Facilities asked about he quantity of isolation valves on the project and TTC said we have one at every unit and a manually s/d on the control valves.

05/15/07 – AHU-1 shutdown & replacement is scheduled for Thursday 05/17. Everyone seems to be on board with the switchover on that day. TTC will confirm with Triton before final shutdown of HVAC-1. Hose bibb change order was processed PDF

05/14/07 – AHU-2 and AHU-3 have been started and are currently running at the building along with the Chiller. We are planning on turning over AHU-1 this Thursday and starting at 6am with that shutdown and rework. TEI has been on site many times over the last two weeks to facilitate the final switchover and changeout of units. AHU-2 was down for nearly 8 hours as the switch took place. JCI has temporary controls on the building for now.

04/18/07 – TTC distributed last week’s meeting minutes. Tim distributed the schedule for the project. The AHU are being set on the curbs today. AHU-3 will have a field erected curb. Const Mtg is this afternoon. TTC has turned in all paperwork for special inspections. Marty Simon said he would review it today or tomorrow and get back to me on it.

04/11/07 – Concrete pour was coordination for today. Kevin check the site yesterday and it appears to be ready. TTC talked to HamCo Inspections today and they still have not approved the Special Inspection part of the permit app. David Puthoff is faxing a sample over to TTC to mimic. TTC distributed meeting minutes with sketches today. PDF

03/30/07 – Triton reported that the units will be delivered on 04/02 at 9am. TE emailed everyone to let them know. Const Mtg 5 was held last Wednesday. Tim Myers was announced as the new PM. Major timing issues were discussed along with the revised schedule which TE will post on the web soon.

03/26/07 – Triton informed TE today that Brian Pastura is leaving the company on Friday March 30. Tim Myers is taking over the PM for Brian on this project. Tim reported that the AHU’s have shipped out of Mexico and should arrive on site this week or next Monday at the latest. TTC will check the schedule for correctness on this situation. TTC contacted Triton and GB Houliston today again about finalizing a change order for the humidifier and getting this piece of equipment on order. TE wants to do this immediately.

03/21/07 – Triton and EAP emailed today discussing RF-1 actuator voltage and damper size. 24 volts and 30×30 was the consensus. Triton to verify. TTC owes 2 const mtg notes to the team and Kevin an explanation about the 2 week scheduling delay in the project. TTC negotiated a revised price from Houliston on the Dri-Steem humidifier. TTC is awaiting confirmation from Triton on the order.

02/19/07 – Last week TEI finished the air device submittal and the EF submittal and sent them back to Triton for ordering. Triton provided an update for the demolition HVAC and plumbing work being completed last week. Tim is now beginning HVAC demo and the duct cleaning is complete in the lower level of the building.

02/02/07 – TEI requested seismic calculations again from Triton (they were due in Feb). Kevin gave Cousins pay application to TTC to approve and TTC sent out an email summary of his field trip with Triton on 02/01/07:

Gentlemen,

My field trip with Triton to discuss ductwork modifications and submittal status revealed several items to share with you. Check my follow up email to this because I want to schedule a regular Const Mtg #04 in a couple of weeks from now. Last Wednesday did not end up as a mtg, sorry for the confusion….

1.) I discussed ductwork changes on AHU-2 (the outdoor 911 unit) and the new layout is going to reduce the downtown when we switch units which is a wonderful thing. We review the unit location on the pad in regards to the interior location and penetration of the wall and distance from the new fence that will eventually be in place. I am really happy with the new layout and approved Triton revised approach for several reasons. 1.) The ductwork routing is simplified. 2.) The downtown between unit changeover is reduced. 3.) The hole in the side of the building is no longer required. 4.) The new AHU-2 can be online while we demolish HVAC-2 and that doesn’t have to happen simultaneously. We are proceeding with the order of this Carrier Unit.

2.) I gave Triton the okay to start limited demolition on HV-1 and HVAC-1 in the basement to the extent that it won’t keep the unit from working but to meet my schedule, they have to start seem demolition of the return ductwork and HV-1 unit.

3.) We have asked for a change order price to raise the AHU-5 condenser up to our new pad level from Triton and will take care of this inside our project budget if the price is acceptable.

4.) The duct cleaning company is scheduled to begin some work in the admin section on 02/12/07. I have asked for this work to be done after hours as the staff is gone at that time and we can judge how loud this cleaning process is going to be before we attempt it in the 911 center.

5.) I am discussing access to the mezzanine with spencer johnson of Cole Russell as I may or may not modify the access door in this project when we are done. I will coordinate that with my architect and the facility safety director (please have steve sears send me the verbiage of what is required IF I decide to put a man door back in the breakroom).

01/17/07 – Chiller sub is 50% reviewed and looks correct. Refrigerant is not as specified in specs.

01/15/07 – After speaking with Majid and Brian at last Friday’s Const Mtg, Triton has provided Carrier Submittals via email for this project. York has lodged a complaint into my voicemail about not being selected and TE has scheduled a meeting with Christian at TE for Wed. TTC will print out the submittals and begin reviewing them immediately.

01/12/07 – Const Mtg with Triton. Concern about equipment and schedule discussed

01/10/07 – Const Mtg was rescheduled to 01/12/06. TTC sent out Const Mtg 01 notes today PDF. Triton called today to report that York equipment is 14 weeks out for delivery.

01/09/07 – IBI permit was not enough for inspection. Banta is taking revised permit app down to IBI to schedule the underslab. Triton met with Banta and confirmed conduit locations up to equipment so that flex tight could be route to the final control locations. Pad pour is scheduled for Thursday/Friday.

01/08/07 – Submittals arrived at 4:45 p.m. last Friday and were logged into the TE system. TTC & TJ are reviewing. The York unit for AHU-2 is not configured correctly and would normally be rejected but we don’t have time for that on this project. TTC to discuss with KM. TEI has faxed IBI dwg review form to IBI for Comm Ctr review for underslab inspection for pad being poured by Cousins.

01/03/07 – Const Mtg 02 was held. Pad pouring were the order of business of the day along with finalizing the schedule. Triton did not have the schedule ready for approval. TTC was frustrated and gave Triton one more week to organize the schedule and ask Triton to take ownership of the project and get submittals into TEI for review.

12/26/06 – TTC and Triton discussed York submittals today. Grady has reported that the factory has already slotted our entire order and that submittals will be here the first week of Jan 07. TEI will review them immediately.

12/07/06 – TA finished electrical submittals. TTC distributed Pre Const Mtg minutes PDF. Triton has to meet owner and TEI on site to discuss pipe routes. Cousin is to provide a change order cost for the fencing on the project.

12/06/06 -Pre Construction meeting was held. TTC to type up notes. Electrical submittals were given to TA. Options for piping was discussed with Facilities about routing more through the building. Building Inspections provided a letter of additional requirements for project.

09/16/06 – TEI is set to have a review set ready for HamCo to review on Wed 09/20. TTC to deliver to County with specs and cost estimates. We are scheduling a review meeting on September 26th.